OpenStudy (anonymous):

The area of a trapezoid is 69 inches squared. Its bases have lengths 9 inches and 14 inches. Find the height of the trapezoid.

OpenStudy (anonymous):

\[area=\frac{ 14+9 }{ 2 }*h=69\] find h.

OpenStudy (marinos):

The area of a trapezoid is given by the formula \[A=\frac{ \left( b_{1}+b_{2} \right)h }{ 2}\] where \[A\] is the area, \[b_{1},b_{2}\] are the bases and \[h\] is the height. Therefore, \[69=\frac{ \left( 9+14 \right)h }{ 2 }\] and so \[h=6\] inches.

OpenStudy (anonymous):

How did you find the Height though?

OpenStudy (mertsj):

What is the formula for the area of a trapezoid?

OpenStudy (anonymous):

\[A=1/2h(b1+b2)\]

OpenStudy (marinos):

@valiant1 We have \[69=\frac{ \left( 9+14 \right)h }{ 2 }\] which implies \[69=\frac{ 23h }{ 2 }\] so \[h=\frac{ 2 }{ 23 }\times69 =6\] Hope this answers your question.

OpenStudy (anonymous):

Aww ok. and one more question how did you know to flip it?